OpenStudy (anonymous):

MEDALS N BECOMING A FAN 1.)Three resistors 2.5 Ω, 3.5 Ω and 4 Ω are connected in parallel. What is the equivalent resistance? 0.94 Ω 9.4 Ω 1.1 Ω 10 Ω 2.) How long would it take for 3.5 C of charge to pass through a cross-sectional area of a wire in which a current of 5 mA passes? 7 s 70 s 700 s 1428 s

OpenStudy (anonymous):

@SolomonZelman @DanJS @Astrophysics

OpenStudy (danjs):

been a minute since physics 2, but i think it is 1/R1 + 1/R2 + 1/R3 maybe

OpenStudy (anonymous):

0.94 Ω

OpenStudy (danjs):

\[\frac{ 1 }{ R _{Tot} } = \frac{ 1 }{ R _{1} }+ \frac{ 1 }{ R _{2} }+\frac{ 1 }{ R _{3} }\]

OpenStudy (danjs):

yeah, just make it 1/0.94 to solve for Rtot

OpenStudy (anonymous):

what about #2 thats the one i was struggling on